OpenStudy (anonymous):

1. Enrique competed in a 47-mile bicycle race last year. He completed the race in 2 hours 34 minutes. This year he would like to complete the race in 2 hours 30 minutes. How many miles per hour should Enrique ride to complete the race in 2 hours 30 minutes? (Use the formula D = rt) (1 point) 18.8 mph 21.3 mph 16.8 mph 23.5 mph

OpenStudy (anonymous):

D = 47 R = ? T = 2.5 you are looking for r, so lets change the equation.. D = rt ---> r = D/T now lets sub in known values... r = 47/2.5 (now divide) r = 18.8 mph <--answer :)